Opened 7 years ago

Last modified 5 years ago

#375 assigned flight processing

UR10/01, flight day 258/2010, Iceland

Reported by: anch Owned by: knpa
Priority: alpha 5 Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by knpa)

Data location: /data/okhotsk1/scratch/arsf/archive-overflow/UR10_01-2010_258_Iceland

Data arrived from ARSF via SATA disk DAPHNE 03 on 7/10/2010

Scientific objective: Not known

Priority: A5

PI: M. Kirkbride

Sensors:

  • Eagle: Delivered 30/11/2010
  • Hawk Delivered 25/02/11
  • Leica LIDAR: Delivered 30/11/2010
  • Digital Photography: Delivered 30/11/2010

Attachments (1)

2582010.sync (478 bytes) - added by anch 7 years ago.
SCT Offsets

Download all attachments as: .zip

Change History (35)

comment:1 Changed 7 years ago by anch

This data is currently held in archive overflow and will need to be moved as space becomes available

comment:2 Changed 7 years ago by anch

  • Summary changed from UR10/01, flight day 258, Iceland to UR10/01, flight day 258/2010, Iceland

comment:3 Changed 7 years ago by iopa

Taking camera.

comment:4 Changed 7 years ago by anch

Applanix and IPAS navigation processing have been done for this project.

Taking hyperspectral.

comment:5 Changed 7 years ago by knpa

Appears to be an issue with the bandset labels in the header files; they start at 900nm which is lower than it should be. Possibly incorrect bandset selected or none specified. Will leave this until end of the year and new Hawk calibration.

comment:6 Changed 7 years ago by iopa

Camera is ready for checking.

Images removed as part of initial testing(.raw);
1-11
Images removed due to cloud/over-exposure/partial images(.raw);
211,297,424-440,473,474,524-527,539,758-762,781-783,787-790

comment:7 Changed 7 years ago by knpa

  • Owner set to anch
  • Status changed from new to assigned

Doing Camera delivery check.

Changed 7 years ago by anch

SCT Offsets

comment:8 Changed 7 years ago by anch

Hyperspectral ready for delivery check.

comment:9 Changed 7 years ago by knpa

Camera delivery check complete. Everything looks fine but PI in readme is C. Cockell which disagrees with this ticket and status spreadsheet (M Kirkbride) . Need to confirm correct PI, check application form and gloucester.

comment:10 Changed 7 years ago by knpa

Beginning Eagle delivery check.

comment:11 Changed 7 years ago by knpa

Delivery check completed.

All looks fine apart from following points:

  • dem is just a symbolic link, need to copy actual file
  • dem name in example azgcorr command should match name of dem file
  • Shouldn't be any misc directory as geoid-spheroid is not needed for ASTER dems. Should also remove mention of this from readme contents and example azgcorr command.
  • Significant cloud in a number of flightlines, should probably mention this in the data quality part of readme

comment:12 Changed 7 years ago by anch

Errors with delivery have been corrected

comment:13 Changed 7 years ago by anch

LiDAR has some gaps at high altitude (Possibly due to cloud/steam over the caldera).

I'm waiting on the results of Kevin's investigation into the same problem on flight day 259b before continuing the classification.

comment:14 Changed 7 years ago by knpa

Phone call from Phil:

This is an issue with the range gate; the plane was flying lower than intended relative to the ground due to an out of date DTM (the volcano had caused the terrain to rise).

Can't do anything about these gaps, can resume processing.

comment:15 Changed 7 years ago by anch

  • Description modified (diff)

Eagle, LIDAR, and camera delivered

comment:16 Changed 7 years ago by anch

Starting Hawk processing.

comment:17 Changed 7 years ago by anch

Hawk submitted for delivery check

comment:18 Changed 6 years ago by emca

Delivery checked Hawk data at
~airborne/workspace/UR10_01-2010_258_Iceland/delivery/20110111

Data all looks good.
Only thing is there is a flight letter on the logsheet (258a) but doesn't appear anywhere else so may just be a mistake on the logsheet. Otherwise ready to deliver.

comment:19 Changed 6 years ago by knpa

Preparing Hawk to be delivered, putting on disk with HS from 259b

comment:20 Changed 6 years ago by knpa

  • Description modified (diff)

Hawk delivered to M. Kirkbride at Dundee (25/02/11)

comment:21 Changed 5 years ago by knpa

Reprocessing HS with APL.

comment:22 Changed 5 years ago by knpa

Stage 2 completed but line 22 E+H missing - APL claims they are not covered by nav file.

comment:23 Changed 5 years ago by knpa

Fixed problem.

Re-running stage 2 as things got a bit confused.

Just for Eagle at this stage - still waiting for Hawk bad pixel stuff.

comment:24 Changed 5 years ago by knpa

Re-running stage 2 (E+H) as apl chain has been updated to generate additional masked files.

comment:25 Changed 5 years ago by knpa

Re-ran stage 2 for Hawk as aplcal has been updated in relation to the bad pixel file.

comment:26 Changed 5 years ago by knpa

  • Milestone set to 2010 data processing completion
  • Owner changed from anch to knpa

comment:27 Changed 5 years ago by knpa

Stage 2 complete.

Had to change a couple of SCTs.

New correct values:

h4 -0.06
h22 1.58

comment:28 Changed 5 years ago by adbe

Hyperspectral reprocessing delivery check:

  • Should the bandlist for the example commands be ALL
  • Screenshots are at the end of the ReadMe (remove [htb])
  • Mention dropped scans in ReadMe

comment:29 Changed 5 years ago by knpa

Moved screenshots to correct place in ReadMe.

Don't think example command bandlist should be ALL as fully mapped files are already provided.

There are no regular dropped scan lines so no need to mention in ReadMe.

PI has not requested that this project be delivered so reprocessing is finished.

comment:30 Changed 5 years ago by knpa

  • Component changed from Processing: general to Archiving

Beginning archiving (of original processing)

comment:31 Changed 5 years ago by knpa

r-syncing from workspace to repository.

comment:32 Changed 5 years ago by knpa

There is tabi data present.

comment:33 Changed 5 years ago by knpa

Finished preparing for archiving, made available to NEODC.

comment:34 Changed 5 years ago by knpa

This was backed-up by NEODC and moved to archive disk a couple of months ago.

Deleting workspace version (original processing). Leaving ticket open for archiving reprocessing.

Note: See TracTickets for help on using tickets.